OpenStudy (barrelracing):

Solve 3x2 + 6x + 15 = 0 −1 ± i − 1 ± 2i −2 ± i −2 ± 2i

OpenStudy (barrelracing):

@freckles

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle x^2+2x+1=(x+1)^2 }\) \(\large\color{black}{ \displaystyle 3(x^2+2x+1)=3x^2+6x+9=3(x+1)^2 }\) *~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~* \(\large\color{black}{ \displaystyle 3x^2+6x+15=0 }\) \(\large\color{black}{ \displaystyle 3x^2+6x+9+6=0 }\) \(\large\color{black}{ \displaystyle 3x^2+6x+9=-6 }\) \(\large\color{black}{ \displaystyle 3(x+1)^2=-6 }\) \(\large\color{black}{ \displaystyle (x+1)^2=-2 }\)

OpenStudy (solomonzelman):

and then proceed by taking the square root of both sides and solving for x.

OpenStudy (barrelracing):

so c?